Femtosecond Photoassociation Spectroscopy (FPAS)

\underline{Una Marvet}, Marcos Dantus, Department of Chemistry, Michigan State University, East Lansing, MI.

The method femtosecond photoassociation spectroscopy (FPAS) is applied to the association of gas-phase mercury atoms to form dimers in the D1u state. Coherent bond formation is demonstrated by rotational anisotropy in the product.